There are 60 science students tn a



Secondary School. 35 of whom study



Chemistry and 30 study Technical



Drawing, 12 out of those students atudy



Biology and chemistry but not Technical



Drawing, 10 study Chemistry but



neither biology nor Technical Drawing 11




study only Technical Drawing and



neither Biology nor Chemistry, 10 also



study Chemistry and Technical Drawing



only




a. How many mdents study all the three



subjects?



b. How many students study biology and



Technical Drawing but not Chemistry?



C. How many students study Biology only?



d How many students study biology all



together.

1
Expert's answer
2022-03-30T02:52:22-0400

"\varOmega -all\,\,students\\C-chemistry\\T-technical\,\,Drawing\\B-bio\log y\\\left| \varOmega \right|=60,\left| C \right|=35,\left| T \right|=30,\left| C\cap \bar{T}\cap B \right|=12,\left| C\cap \bar{T}\cap \bar{B} \right|=10,\left| \bar{C}\cap T\cap \bar{B} \right|=11,\left| C\cap T\cap \bar{B} \right|=10\\a:\\\left| C\cap T\cap B \right|=\left| C \right|-\left| C\cap \bar{T}\cap \bar{B} \right|-\left| C\cap \bar{T}\cap B \right|-\left| C\cap T\cap \bar{B} \right|=35-10-12-10=3\\b:\\\left| \bar{C}\cap T\cap B \right|=\left| T \right|-\left| C\cap T\cap \bar{B} \right|-\left| C\cap T\cap B \right|-\left| \bar{C}\cap T\cap \bar{B} \right|=30-10-3-11=6\\c:\\\left| \bar{C}\cap \bar{T}\cap B \right|=\left| \varOmega \right|-\left| C\cap T\cap B \right|-\left| C\cap \bar{T}\cap B \right|-\left| \bar{C}\cap T\cap B \right|-\left| C\cap \bar{T}\cap \bar{B} \right|-\left| C\cap T\cap \bar{B} \right|-\left| \bar{C}\cap T\cap \bar{B} \right|=\\=60-3-12-6-10-10-11=8\\d:\\\left| B \right|=\left| \bar{C}\cap \bar{T}\cap B \right|+\left| \bar{C}\cap T\cap B \right|+\left| C\cap \bar{T}\cap B \right|+\left| C\cap T\cap B \right|=8+6+12+3=29"